OpenStudy (anonymous):

help! George runs an outlet where he sells personalized T-shirts. He spends $35 to rent the printing machine and $25 for phone service each month. If each T-shirt costs $5.80, and he sells them for $10, George will start making a profit after selling shirts ? in a month.

OpenStudy (sepeario):

Okay let's do some math here. \[35+25+5.8x=E\], where x is the number of shirts and E is expenditure. Then the equation for revenue is obviously \[10x\] so the inequality is \[70+5.8x <10x\]

OpenStudy (sepeario):

For the profit to be positive.

OpenStudy (anonymous):

my choices were 4, 14, 15, and 42

OpenStudy (anonymous):

I would simply do it by finding the intersection of two lines. One line: your average cost to make a shirt (70+5.8x) Second line: you profit (10x) Set them equal to each other and solve

OpenStudy (sepeario):

pellet I stuffed up :( it's supposed to be 60 not 70. Sorry. \[4.2x >60, x >14\]

OpenStudy (anonymous):

So 70+5.8x=10x solve I think that works?

OpenStudy (sepeario):

That's essentially what I did.

OpenStudy (sepeario):

In fact that's exactly what I did.

OpenStudy (anonymous):

16.6667 but thats not right

OpenStudy (sepeario):

No read my next comment it's 14

OpenStudy (sepeario):

Above

OpenStudy (anonymous):

(35+25=60 not 70) we both messed up at first binks

OpenStudy (anonymous):

(so in my equation replace 70 with a 60

OpenStudy (anonymous):

And yes sepeario I know you were correct. I was only simplifying what you said

OpenStudy (anonymous):

Does this make sense binks?

OpenStudy (anonymous):

thanks sorry i had to go Thanks for the help!

OpenStudy (anonymous):

wish I could give you both medals thanks so much!!